# Evaluating LLMs’ Mathematical and Coding Competency through Ontology-guided Interventions

### v4

@@ I.3 Representational Perturbation – Format Change
Alternative Format:
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] Maria has 4 cuboids. Each cuboid is 7 feet by 9 feet by 6 feet. The walls are 2 feet thick. What is the total volume of all the cuboids? Realworld Usecase: Frame the requirement in docstring into a problem that will happen in a realworld scenario. Changed from G.1:
@@
Change Sequence: Change the order of the information and names of the variables that is originally presented in the question. Change Sequence: Change the sequence of information given in the question without affecting the solvability of the question. Changed from F.2: {mdframed}[backgroundcolor=pink!20] The walls of John’s boxes are 1 inch thick. Each of these boxes measures 5 inches by 6 inches by 4 inches. John has 3 boxes. What is the total inner volume of all 3 boxes? Parameter Sequence: Change the sequence of the input parameter and change the input parameter names. Changed from G.2
@@
Close Format: Rewrite the sentence as a closed-format question that evaluates the correctness of possible answers. True False: Evaluate a potentially misleading answer and check the correctness of the answer. Changed from F.1: {mdframed}[backgroundcolor=pink!20] A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth $5,000 or electronic gadgets worth $8,000. His financial advisor speculates that the jewelry market will go up 2.5% while the electronic gadgets market will rise 1.2% within the same month. If the merchant is looking to maximize profit at the end of this month by making a choice, how much profit would this be? Evaluate the correctness of this answer with respect to the above question: $96. True False: Check if a given code answer can solve the requirement in docstring. Changed from G.2: Evaluate whether the solution below is the correct solution for the coding question, True or False?
@@
Data Restructuring: Change the layout, organization of the data presented in the question. Value Structuring: Arrange the variables inside the question in a tabular format. Changed from F.2:
| Variable | Value | |----------|-------| | a | 3 | | b | 5 | | c | 6 | | d | 4 | | e | 1 |
@@
John has ‘a’ boxes. Each box is ‘b’ inches by ‘c’ inches by ‘d’ inches in dimensions. The walls are ‘e’ inch thick. What is the total inner volume of all the ‘a’ boxes? Complex Docstring: Elaborate the documentation string by exhaustively detailing more conditional pathway within the code. Changed from G.1:
@@
Identical Problem: Check if the two question or code are identical in describing or solving the same problem. Identical Question: If two questions requires exactly the same framework or thinking procedure to solve. Changed from F.1: {mdframed}[backgroundcolor=pink!20] Question 1: A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth $5,000 or electronic gadgets worth $8,000. His financial advisor speculates that the jewelry market will go up 2.5% while the electronic gadgets market will rise 1.2% within the same month. If the merchant is looking to maximize profit at the end of this month by making a choice, how much profit would this be? Question 2: An investor is unsure of which investment to make: gold valued at $10,000 or stocks valued at $15,000. His financial consultant predicts that the gold market will inflate by 3.5% while the stock market will increase by 2.2% over the next quarter. If the investor wants to achieve the highest return on his investment at the end of this quarter, how much would his initial investment be? Does Question 1 and Question 2 require identical steps to answer? Identical Code: Are the two solutions to the question identical in terms of their functionality? Changed from G.3 Is function_1 and function_2 identical in terms of its functionality?
@@
Reasoning Format: The format for the final answer should be converted in a certain way
@@
Binary Coded: Answer the final question in base-n. Changed from F.2: {mdframed}[backgroundcolor=pink!20] Answer the following question with only base-2 coded values. Question: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? No Keyword: The solution should not use a specific python keyword. For example, “for” or “while”. Changed from G.2: Answer the coding function below without using python keywords: "while", "for" in the solution
@@
Reasoning Style: The reasoning steps should be performed in a certain style. X Language (Math): Give the answer in certain language from Spanish, Chinese, Bengali, English, French Changed from F.2: Answer the following question with only Chinese language, because I do not understand English. {mdframed}[backgroundcolor=pink!20] Question: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? X Language (Code): Give the code answer in another coding language. Changed from G.1: Answer the coding question below;
@@
Alternative Answer: Find the alternative solutions to existing solution. Alternative Answer (Math) : Give an alternative solution that is different from the standard reasoning steps, but arrives at the same correct final answer. Changed from F.2: {mdframed}[backgroundcolor=pink!20] Give an different step-by-step solution to calculate the answer to the following question. Make sure the solution is different from the solution below. Question: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? Solution: The walls subtract a (1 + 1) = 2 inches from each dimension. So, each box has a reduced width of (5 - 2) = 3 inches, reduced length of (6 - 2) = 4 inches and reduced height of (4 - 2) = 2 inches. So the inner volume of each box is 3 * 4 * 2 = 24 cubic inches. The total inner volume of 3 boxes are 3 * 24 = 72 cubic inches. Alternative Step by Step Solution: Alternative Answer (Code): Find an alternative solution to existing coding solution. Changed from G.1:
@@
New Rule: Integrate a new rule into the original question that requires the solution follow the new rule. This type tests the model’s ability to adapt to new ruels and knowledge and use it inside the solution. Define Rules:Define a new mathematical rule that will change how the toolbox (commonsense knowledge) is applied during calculation. Changed from F.2: {mdframed}[backgroundcolor=pink!20] In a parallel universe, John has 3 boxes. Each box has peculiar dimensions: 5 quarks by 6 quarks by 4 quarks with walls that are 1 quark thick. In this universe, the total inner volume of a box is calculated by using the Illusory Volume operation, represented as IV. The IV operation is defined as: (length * width * height) - (number_of_walls * thickness_of_each_wall). What is the total inner volume of all 3 boxes? Simple Name: The generated code should only have variables names in a certain format. Changed from G.1: Answer the coding question below and only use 6 letter word for each variable names inside the solution
@@
Self Consistency Prompting: We randomly picked one question answer pair in the same category from our maths dataset More and prepend it to the front of the perturbed question as a one shot demonstration. Then we use the below prompt template for Self Consistency prompting. We sample the generation three times and get the final answer by majority voting. In case of tie, we randomly pick an answer. {mdframed}[backgroundcolor=gray!10] Given the oneshot demonstration of a question and its final answer, Solve the #question# step by step before giving the final answer. Do not directly give the final answer.
#Demonstration Question#: demonstration question #Demonstration Final Answer#: demonstration answer
@@
Program of Thought Prompting: We use the following prompt template for the program of thought experiments: {mdframed}[backgroundcolor=gray!10] Instruction: You are an experienced professional skilled in using python programs to solve math related problems. Solve the question below using python programs, You will only write code blocks. Problem: Question
@@
The detailed results across the perturbation categories for all the models are illustrated in Tables 8 and 9.
@@
Question Simplification | Remove Constraint | 5 | 5 | 5 | 5 | 4 |
@@
Dimension | Category | GPT-4 | ChatGPT | Gemini | CodeLlama | Llama2-Chat |
@@
Question Simplification | Remove Constraint | 4 | 4 | 4 | 4 | 2 |

### v5

@@ I.3 Question Format
Format Change:
G36.
Setting Rephrase: Rephrase the question in another setting.
Change Setting: Rephrase by changing the application setting and values inside the information, while keeping the core mathematical structure intact.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] Maria has 4 cuboids. Each cuboid is 7 feet by 9 feet by 6 feet. The walls are 2 feet thick. What is the total volume of all the cuboids? Realworld Usecase: Frame the requirement in docstring into a problem that will happen in a realworld scenario. Changed from G.1:1def switch_text_case(text: str) -> str:2 """3 Imagine you’re working on a document and you’ve mistaken the case in the text you write. You wrote all the lower case letters in uppercase and vice versa, suppose you want to correct all of them using python.4 """ -
G37.
Change Sequence: Change the order of the information and names of the variables that is originally presented in the question.
Change Sequence: Change the sequence of information given in the question without affecting the solvability of the question.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] The walls of John’s boxes are 1 inch thick. Each of these boxes measures 5 inches by 6 inches by 4 inches. John has 3 boxes. What is the total inner volume of all 3 boxes? Parameter Sequence: Change the sequence of the input parameter and change the input parameter names.
Changed from G.21def munchee_bunchee(xray: int, yoyo: int) -> int:23 """ Return a common divisor that is the largest of two integers xray and yoyo4 >>> munchee_bunchee(3, 5)5 16 >>> munchee_bunchee(25, 15)7 58 """ -
G38.
Close Format: Rewrite the sentence as a closed-format question that evaluates the correctness of possible answers.
True False: Evaluate a potentially misleading answer and check the correctness of the answer.
@@
Changed from F.1: {mdframed}[backgroundcolor=pink!20] A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth $5,000 or electronic gadgets worth $8,000. His financial advisor speculates that the jewelry market will go up 2.5% while the electronic gadgets market will rise 1.2% within the same month. If the merchant is looking to maximize profit at the end of this month by making a choice, how much profit would this be? Evaluate the correctness of this answer with respect to the above question: $96. True False: Check if a given code answer can solve the requirement in docstring. Changed from G.2: Evaluate whether the solution below is the correct solution for the coding question, True or False?1Function:23def greatest_common_divisor(a: int, b: int) -> int:45 """ Return a greatest common divisor of two integers a and b6 >>> greatest_common_divisor(3, 5)7 18 >>> greatest_common_divisor(25, 15)9 510 """111213Solution:1415 while a:16 a, b = a % b, a17 return b -
G39.
Data Restructuring: Change the layout, organization of the data presented in the question.
@@
Value Structuring: Arrange the variables inside the question in a tabular format.
Changed from F.2:| Variable | Value | |----------|-------| | a | 3 | | b | 5 | | c | 6 | | d | 4 | | e | 1 |
@@
John has ‘a’ boxes. Each box is ‘b’ inches by ‘c’ inches by ‘d’ inches in dimensions. The walls are ‘e’ inch thick. What is the total inner volume of all the ‘a’ boxes?
Complex Docstring: Elaborate the documentation string by exhaustively detailing more conditional pathway within the code.
Changed from G.1:1def function(string: str = None) -> str:2 """3 For any specified sequence of alphabetical characters, interspersed with spaces, numerical digits, and various symbols, implement a sophisticated transformation algorithm designed to selectively convert each alphabetical character from its current case representation, either lowercase or uppercase, to its diametrically opposite case representation. This algorithm ensures that every character initially presented in lowercase is meticulously transmuted to uppercase, and conversely, every character originally in uppercase is converted to lowercase, while meticulously preserving the integrity and original positioning of spaces, numerical digits, and any other non-alphabetical symbols, leaving these elements unaltered within the sequence.4 >>> function(’Hello’)5 ’hELLO’6 """
G36.
Format Comparison:
G40.
Identical Problem: Check if the two question or code are identical in describing or solving the same problem.
Identical Question: If two questions requires exactly the same framework or thinking procedure to solve.
@@
Changed from F.1: {mdframed}[backgroundcolor=pink!20] Question 1: A merchant wants to make a choice of purchase between 2 purchase plans: jewelry worth $5,000 or electronic gadgets worth $8,000. His financial advisor speculates that the jewelry market will go up 2.5% while the electronic gadgets market will rise 1.2% within the same month. If the merchant is looking to maximize profit at the end of this month by making a choice, how much profit would this be? Question 2: An investor is unsure of which investment to make: gold valued at $10,000 or stocks valued at $15,000. His financial consultant predicts that the gold market will inflate by 3.5% while the stock market will increase by 2.2% over the next quarter. If the investor wants to achieve the highest return on his investment at the end of this quarter, how much would his initial investment be? Does Question 1 and Question 2 require identical steps to answer? Identical Code: Are the two solutions to the question identical in terms of their functionality?
Changed from G.3 Is function_1 and function_2 identical in terms of its functionality?1Code 1:2def function(xs: list):3 return [(i * x) for i, x in enumerate(xs)][1:]4Code 2:5def function(xs: list):6 derivative = [i * xs[i] for i in range(1, len(xs))]
G40.
I.4 Answer Format
@@
Reasoning Format: The format for the final answer should be converted in a certain way.
Binary Coded: Answer the final question in base-n.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] Answer the following question with only base-2 coded values. Question: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? No Keyword: The solution should not use a specific python keyword. For example, “for” or “while”. Changed from G.2: Answer the coding function below without using python keywords: "while", "for" in the solution1def greatest_common_divisor(a: int, b: int) -> int:23 """ Return a greatest common divisor of two integers a and b4 >>> greatest_common_divisor(3, 5)5 16 >>> greatest_common_divisor(25, 15)7 58 """ -
G42.
Reasoning Style: The reasoning steps should be performed in a certain style.
X Language (Math): Give the answer in certain language from Spanish, Chinese, Bengali, English, French
Changed from F.2: Answer the following question with only Chinese language, because I do not understand English.
@@ X Language (Math): Give the answer in certain language from Spanish, Chinese, Bengali, English, French
{mdframed}[backgroundcolor=pink!20] Question: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? X Language (Code): Give the code answer in another coding language.
Changed from G.1: Answer the coding question below;1func flipCase(str string) string {2// flipCase takes a string and flips the case of each character: lowercase to uppercase and uppercase to lowercase.34} -
G43.
Alternative Answer: Find the alternative solutions to existing solution.
Alternative Answer (Math) : Give an alternative solution that is different from the standard reasoning steps, but arrives at the same correct final answer.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] Give an different step-by-step solution to calculate the answer to the following question. Make sure the solution is different from the solution below. Question: John has 3 boxes. Each box is 5 inches by 6 inches by 4 inches. The walls are 1 inch thick. What is the total inner volume of all 3 boxes? Solution: The walls subtract a (1 + 1) = 2 inches from each dimension. So, each box has a reduced width of (5 - 2) = 3 inches, reduced length of (6 - 2) = 4 inches and reduced height of (4 - 2) = 2 inches. So the inner volume of each box is 3 * 4 * 2 = 24 cubic inches. The total inner volume of 3 boxes are 3 * 24 = 72 cubic inches. Alternative Step by Step Solution: Alternative Answer (Code): Find an alternative solution to existing coding solution. Changed from G.1:1Find a different solution other than:2def flip_case(string: str) -> str:34 return string.swapcase() -
G44.
New Rule: Integrate a new rule into the original question that requires the solution follow the new rule. This type tests the model’s ability to adapt to new ruels and knowledge and use it inside the solution.
Define Rules:Define a new mathematical rule that will change how the toolbox (commonsense knowledge) is applied during calculation.
@@
Changed from F.2: {mdframed}[backgroundcolor=pink!20] In a parallel universe, John has 3 boxes. Each box has peculiar dimensions: 5 quarks by 6 quarks by 4 quarks with walls that are 1 quark thick. In this universe, the total inner volume of a box is calculated by using the Illusory Volume operation, represented as IV. The IV operation is defined as: (length * width * height) - (number_of_walls * thickness_of_each_wall). What is the total inner volume of all 3 boxes? Simple Name: The generated code should only have variables names in a certain format.
Changed from G.1: Answer the coding question below and only use 6 letter word for each variable names inside the solution1def flip_case(string: str) -> str:2 """For a given string, flip lowercase characters to uppercase and uppercase to lowercase.3 >>> flip_case(’Hello’)4 ’hELLO’5 """
G41.
@@
Self Consistency Prompting:
We randomly picked one question answer pair in the same category from our math dataset GSMore and prepend it to the front of the perturbed question as a one shot demonstration. Then we use the below prompt template for Self Consistency prompting. We sample the generation three times and get the final answer by majority voting. In case of tie, we randomly pick an answer.
{mdframed}[backgroundcolor=gray!10]
Given the oneshot demonstration of a question and its final answer, Solve the #question# step by step before giving the final answer. Do not directly give the final answer.
#Demonstration Question#: demonstration question
#Demonstration Final Answer#: demonstration answer
@@
Program of Thought Prompting:
We use the following prompt template for the program of thought experiments:
{mdframed}[backgroundcolor=gray!10]
Instruction: You are an experienced professional skilled in using python programs to solve math related problems. Solve the question below using python programs, You will only write code blocks.
Problem: Question
@@
The detailed results across the perturbation categories for all the models are illustrated in Tables 9 and 10.
@@
| Granularity Adjustment | Remove Constraint | 5 | 5 | 5 | 5 | 4 |
@@
| Dimension | Category | GPT-4 | GPT-3.5 | Gemini | CodeLlama | Llama2-Chat |
@@
| Granularity Adjustment | Remove Constraint | 4 | 4 | 4 | 4 | 2 |